OpenStudy (anonymous):

PLEASE HELP ME What is the value for the x in the proportion x + 8 over 5x - 2 equals 3 over 8?

OpenStudy (anonymous):

\[x + 8/5x-2 = 3/8\]

OpenStudy (anonymous):

Cross multiply then solve for x, so \[8\times8=64 \] \[8+x=8x\] \[(5x)\times3=15x\] \[3\times-2=-6\] so it would be \[8x+64=15x-6\] then 70=7x so x=10

OpenStudy (anonymous):

\[\frac{x+8}{5x-2}=\frac{3}{8}\] @Ashley1997 please don't do the work for them, that's not what the site is here for. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

I know. That’s why I did the work first before the answer because they would have to see the work to get the answer. I think it's better if I do the work in the hopes that they might read it instead of just giving them the answer.